At a certain temperature, the Kp for the decomposition of H2S is 0.719.

H2S(g)−⇀H2(g)+S(g)

Initially, only H2S is present at a pressure of 0.193 atm in a closed container. What is the total pressure in the container at equilibrium?
1
Expert's answer
2020-03-31T13:44:32-0400

"H_2S(g)\longrightarrow H_2(g) + S(g)"


here it is given as Kp=0.719 atm and p(H2S)=.193 atm

so we will take x part of H2S is change,

so we can write the equilibrium reaction as

"H_2S(g)\longrightarrow H_2(g) + S(g)"

0.193 0 0

0.193-x x x


we can write for equilibrium constant Kp as


"K_p= \frac{P_{H_2}\times P_{S}}{P_{H_2S}}"


"0.719= \frac{x\times x}{0.193-x}"


0.139-0.719"x=x^2"


"x^2+0.719x-0.139=0"


using quadratic formula we can find


"x=\frac{-0.719 \pm\sqrt{(.719^2-(4\times1\times(-.139)))}}{2}"


x= "\frac{-0.719+1.04}{2}=0.1605,x=\frac{-0.719-1.04}{2}=-0.8795"


so here , "P_{H_2}=P_S=0.1605 atm, and P_{H_2S}=0.719-0.1605=0.5585 atm"


Here total pressure will be sum of all


PT= 0.1605+0.1605+0.5585=0.8795 atm

"P_T= 0.8795 atm"